# Scalable Dynamic Origin-Destination Demand Estimation Enhanced by High-Resolution Satellite Imagery Data

### v1

@@ Abstract
This study presents a novel integrated framework for dynamic origin-destination demand estimation (DODE) in multi-class mesoscopic network models, leveraging high-resolution satellite imagery together with conventional traffic data from local sensors. Unlike sparse local detectors, satellite imagery offers consistent, city-wide road and traffic information of both parking and moving vehicles, overcoming data availability limitations. To extract information from imagery data, we design a computer vision pipeline for class-specific vehicle detection and map matching, generating link-level traffic density observations by vehicle class. Building upon this information, we formulate a computational graph-based DODE model that calibrates dynamic network states by jointly matching observed traffic counts and travel times from local sensors with density measurements derived from satellite imagery. To assess the accuracy and scalability of the proposed framework, we conduct a series of numerical experiments using both synthetic and real-world data. The results of out-of-sample tests demonstrate that supplementing traditional data with satellite-derived density significantly improves estimation performance, especially for links without local sensors. Real-world experiments also confirm the framework’s capability to handle large-scale networks, supporting its potential for practical deployment in cities of varying sizes. Sensitivity analysis further evaluates the impact of data quality related to satellite imagery data.
@@ 1 Introduction
The widespread availability of spatio-temporal data has created new opportunities for advancing computational tools to model network flows, individual traveler behavior, and travel demand in dynamic transportation networks. Recent developments in sensing technologies and artificial intelligence are revolutionizing traditional models, making them more data-driven, scalable, and effective for complex, large-scale networks. Dynamic Origin-destination Demand Estimation (DODE) is a foundational prerequisite for dynamic network models to accurately reproduce the status quo spatio-temporal network conditions, supporting traffic assignment [Pi et al., 2019] and control strategies [Ye et al., 2019, Liu, Ma and Qian, 2023, Ke et al., 2025]. DODE studies can be broadly categorized into model-based methods, which embed physics-informed traffic assignment models, and model-free methods, which formulate the problem using data-driven techniques without traffic assignment constraints. Recent studies have merged the advantages of both approaches, reformulating DODE within learning-based frameworks and demonstrating promise in large-scale applications [Wu et al., 2018, Liu and Yin, 2025]. Despite extensive research over the past few decades, DODE remains challenging because it is inherently a high data-demanding problem due to its high-dimensional and under-deterministic nature. Accurate and reliable DODE, particularly in large-scale networks, requires the integration of diverse and comprehensive data sources.
@@
Existing DODE studies reply on either single or multiple data sources from fixed traffic detectors (e.g., traffic count), mobile sensors (e.g., probe vehicle traces, speeds) and activity-based data (e.g., surveys, mobile device traces). However, one major limitation shared by these data sources is that they capture only a subset of travelers within the system due to the sparsity of sensor geographical coverage, low penetration rate of mobile sensors, low sample rate of survey and privacy issues of data collection. Prior studies demonstrated that the performance of DODE is highly sensitive to sensor placement [Zhu et al., 2016, 2017] while optimal sensor deployment requires long-term planning for new installations and routine maintenance for deployed sensors, both of which can be costly and time-consuming [Zhou and List, 2010, Salari et al., 2021, Li et al., 2023, Hu and Fan, 2024].
@@
In recent years, satellite remote sensing data have emerged promising new data sources for city-wide transportation models, overcoming weakness of localized and mobile ground-based sensors [Sheehan et al., 2023, Gagliardi et al., 2023]. It has demonstrated effectiveness in monitoring both transportation infrastructure conditions and network-wide traffic dynamics. These data often have larger geographical coverage, and provide detailed, consistent information of all individual vehicles detected and road features in city-wide networks. Advances in computer vision and machine learning have significantly enhanced the information extraction process from these satellite images, including vehicle detection [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016] and road segmentation [Dai et al., 2020, Mei et al., 2021]. Fusing this emerging data source, which is unconstrained by spatial locations, into DODE has great potential to make it more scalable, accurate and economic.
@@
This study investigates the potential of incorporating high-resolution imagery data collected by flexible sky-view sensing techniques such as satellite into a computation graph-based DODE framework to enhance model scalability and performance in replicating real-world network-wise traffic conditions. The contributions are summarized as follows:
@@
    We develop a novel dynamic OD demand estimation framework fusing multi-modal data including tabular data from localized detectors and imagery data from satellite. The satellite imagery data which is not restricted by the location, complementing the limitations of local sensors and having potentials for scalable applications.
@@
    We develop a computer vision pipeline to extract information from high-resolution satellite imagery data, including detecting both moving vehicles and on-street parking vehicles by class, matching vehicles to links and computing link-level traffic densities which is a new observation in DODE.
@@
    We design a general mesoscopic dynamic network loading (DNL) model compatible with traffic density estimation, and develop a DODE framework that integrates multi-source traffic condition data collected from local sensors and satellite images. The framework is formulated on computational graphs and solved efficiently on large-scale networks. Additionally, we address the data availability and inconsistency issues of multi-source data fusion.
  • 4.
    We use both synthetic and real-world data on toy and large-scale networks to evaluate model performance improvement when incorporating information from satellite imagery data. We perform sensitivity analysis to examine model robustness with respect to data quality, including computer vision sensing errors and imagery collection frequency.
The organization of the remaining paper is as follows: Section 2 summarizes related studies in DODE and usage of satellite imagery in transportation research. Section 3 introduces the proposed DODE framework integrating multi-source data including satellite imagery information and the general DNL model for density estimation. Section 4 conducts numerical experiments using both real-world multi-source data and synthetic data. Sensitivity analysis is also performed to test model’s robustness to data quality of density information extracted from satellite imagery. Section 5 summarizes findings and provides future research directions.
@@ 2.1 Dynamic OD Demand Estimation
From a modeling perspective, DODE studies can be broadly categorized into model-based, model-free and overlapping approaches. Model-based approaches typically use a bi-level optimization structure, where upper level minimizes observed and estimated traffic conditions, and the lower level adopts a Dynamic Traffic Assignment (DTA) model to capture travel behavior and network traffic dynamics. In contrast, model-free methods do not use physical models and rely on data-driven approaches using various machine learning architectures. Overlapping approaches leverage advantages of both model-based and model-free approaches.
Regardless of the modeling approach, DODE problems are highly data-demanding. The traffic information data used in DODE are collected using either fixed detectors or mobile detectors. Fixed detectors has permanently fixed locations and provide continuous localized traffic conditions at specific locations. Some examples are traffic count data collected by loop detectors, automated vehicle identification (AVI) data collected by communication with the vehicle tags and barrier devices, and license plate recognition (LPR) data capture by fixed point monitoring cameras. Mobile detectors refer to portable device attached to a subset of travelers in the system and collect successive localized traffic conditions based on location changes of these travelers. Some examples are link speed data collected by probe vehicles on the link, complete vehicle traces of probe vehicles and human’s activity trajectories collected by mobile phone location tracking. Table 1 summarizes representative studies utilizing various data sources and modeling approaches.
@@
Dixon and Rilett [2002] automated vehicle identification &\&& vehicle traces model-free
@@
Ma et al. [2020] traffic count &\&& speed model-based
@@
Cao et al. [2021] automated vehicle identification &\&& vehicle traces model-free
@@
However, both fixed and mobile detectors have limitations. Fixed detector data relies heavily on sensor locations and often sparsely distributed in large networks, making it challenging to infer full network demand [Hu and Fan, 2024]. Additionally, maintaining installed sensors and planning new locations can be costly especially in large-scale networks, limiting scalability and sustainability. Mobile detectors, while more flexible, depends on penetration rates and a subset of travelers may not well represent the entire population [He et al., 2024]. Furthermore, both sensing categories suffer from inconsistent data quality at different times, different locations and different data sources, leading to contradictory conditions collected for the same link. To overcome these limitations, some studies have fused multiple complementary data sources, usually from different detectors, to improve DODE performance. Frederix et al. [2010] combined additional link-level density information into DODE, outperforming traditional estimation approaches using only traffic flow, which is not effective in discerning between congested and uncongested traffic states. Carrese et al. [2017] incorporate probe vehicle traces data with link traffic flow measurements to distinguish congestion states and enhance DODE reliability.
Recent studies have reformulated OD demand estimation and related problems within machine learning framework, proposing computational graph-based approaches that efficiently fuse multiple data sources. Wu et al. [2018] proposed a multi-layered computational graph representation to structurally model different levels of travel demand variables and individual travel behavior parameters. Ma et al. [2020] formulated the DODE problem on a computational graph utilizing multi-source multi-class traffic count and speed data. Lu et al. [2023] proposed an physics-informed framework for joint traffic state and queue profile estimation on layered computational graph by leveraging heterogeneous data sources. Kim et al. [2024] further developed a comprehensive analytical mathematical formulation to integrate transport demand and supply models in a computational graph framework. Guarda et al. [2024] developed a computational graph framework to estimate network flow and travel behavior using day-to-day system-level data and Guarda and Qian [2024] uses this data-driven framework to estimate traffic conditions in unobserved network locations. Liu and Qian [2024] estimated curbside usage demand patterns based on additional curb event monitoring data along with traffic count and speed information. Liu, Yin, Bai and Grimm [2023] and Liu and Yin [2025] embedded neural networks in implicit layers of an end-to-end learning framework to directly learns travel choice preferences, supply and demand components and the equilibrium state from multi-day link flow observations. Du et al. [2025] developed a fully differentiable end-to-end simulation-based optimization framework to calibrate passenger route choice using smart card and train loadings data. These studies demonstrated the potential of integrating diverse data sources, especially new data sources that can well address the concern of inconsistency information, within unified learning frameworks, leading to robust and precise demand estimation.
@@ 2.2 Satellite imagery in transportation research
Advances in computer vision and sensing has unlocked the potential of remote sensing data for enhancing city-wide infrastructure monitoring Gagliardi et al. [2023] and traffic condition monitoring [Sheehan et al., 2023]. High-resolution imagery data captures consistent, city-wide traffic conditions of all links at the same time, overcoming the limitations of ground-based sensors. Some pilot studies prove that the effectiveness of traffic monitoring using satellite imagery data [McCord et al., 2003, Larsen et al., 2009, 2013, Kaack et al., 2019]. Additionally, Unmanned Aerial Vehicles (UAV) can be another similar but more flexible approach to collect precise imagery or video-based traffic data in small sub-regions of city, due to their rapid speed, cost-effectiveness, and extensive field-of-view [Balamuralidhar et al., 2021, Butilă and Boboc, 2022, Bai and Feng, 2025].
@@
Some pilot studies have explored the utilization of these imagery data for transportation research, including road network extraction [Dai et al., 2020, Mei et al., 2021], vehicle detection and classification [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016]. Some studies have focused on traffic density estimation based on imagery or video collected from satellite [Kopsiaftis and Karantzalos, 2015, Sakai et al., 2019]. Reksten and Salberg [2021] proposed a deep learning-based framework to estimate annual average daily traffic (AADT) using very-high-resolution optical remote-sensing imagery with spatially limited ground-based measurements. Ganji et al. [2022] used Google aerial images together with sparse data from traffic counts to predict traffic volumes. Chen et al. [2021] used traffic density detection data to examine changes in spatio-temporal traffic patterns during the COVID 19 pandemic. He et al. [2021] integrated satellite imagery and GPS trajectories to generate high-resolution accident risk maps. Yu et al. [2024] integrated high-resolution satellite imagery with geo-spatial tabular and crash data to analyze factors contributing to freight truck-related crashes. Horsler et al. [2023] incorporated information from satellite imagery to predict regional road transport emissions, highlighting the potential in environmental impact assessments. Our initial work proves the potential of incorporating satellite imagery into mesoscopic network model calibration but the information extraction from imagery data is not vehicle class specific and the parking information is not considered, leading to limited accuracy in large-scale network applications [Liu et al., 2024].
@@
To the best of our knowledge, none of the existing studies incorporated detailed traffic information of both moving and on-street parking vehicles by class extracted from satellite images directly into calibrating dynamic network models and performing DODE. Incorporating satellite imagery has the potential to fill gaps in areas where ground-based sensors are not available, providing a more complete picture of consistent network conditions to enhance DODE performance.

### v2

@@ Abstract
This study presents a novel integrated framework for dynamic origin-destination demand estimation (DODE) in multi-class mesoscopic network models, incorporating high-resolution satellite imagery together with conventional traffic data from local sensors. Unlike sparse local detectors, satellite imagery offers consistent, city-wide road and traffic information of both parking and moving vehicles, overcoming data availability limitations. To extract information from imagery data, we design a computer vision pipeline for class-specific vehicle detection and map matching, generating link-level traffic density observations by vehicle class. Building upon this information, we formulate a computational graph-based DODE framework that calibrates dynamic network states by jointly matching observed traffic counts/speeds from local sensors with density measurements derived from satellite imagery. To assess the accuracy and robustness of the proposed framework, we conduct a series of numerical experiments using both synthetic and real-world data. The results demonstrate that supplementing traditional data with satellite-derived density significantly improves estimation performance, especially for links without local sensors. Real-world experiments also show the framework’s potential for practical deployment on large-scale networks. Sensitivity analysis further evaluates the impact of data quality related to satellite imagery data.
@@ 1 Introduction
The widespread availability of spatio-temporal data has created new opportunities for advancing computational tools to model network flows, individual traveler behavior, and travel demand in dynamic transportation networks. Recent developments in sensing technologies and artificial intelligence are revolutionizing traditional models, making them more data-driven, scalable, and effective for complex, large-scale applications. Dynamic Origin-Destination Demand Estimation (DODE) is a foundational prerequisite for dynamic network models to accurately reproduce the current spatio-temporal network conditions, supporting traffic assignment [Pi et al., 2019] and control strategies [Ye et al., 2019, Liu et al., 2023a, Wu et al., 2024, 2025]. Traditional DODE studies embed physics-based traffic assignment models in a bi-level model structure, considering both traffic flow propagation and travel behavioral dynamics, while emerging studies formulate the problem using purely data-driven models without traffic assignment constraints. Recent studies have merged the advantages of both approaches, reformulating physics-based DODE within learning-based frameworks and demonstrating promise in large-scale applications [Wu et al., 2018, Liu et al., 2023b, Liu and Yin, 2025, Zhou et al., 2025]. Despite extensive research over the past few decades, DODE remains challenging because it is inherently a highly data-demanding problem due to its high-dimensional and under-determined nature. Accurate and reliable DODE, particularly on large-scale networks, requires the integration of diverse and comprehensive data sources.
@@
Existing DODE studies rely on either single or multiple data sources from fixed traffic detectors (e.g., traffic counts), mobile sensors (e.g., probe vehicle traces, speeds) and activity-based data (e.g., surveys, mobile device traces). However, one major limitation shared by these data sources is that they capture only a subset of travelers within the system due to the sparsity of the geographic coverage of sensors, low penetration rates of mobile sensors, low sample rates of surveys and privacy issues in data collection. Prior studies have demonstrated that the performance of DODE is highly sensitive to sensor placement [Zhu et al., 2016, 2017] while optimal sensor deployment requires long-term planning for new installations and routine maintenance for deployed sensors, both of which can be costly and time-consuming [Zhou and List, 2010, Salari et al., 2021, Li et al., 2023, Hu and Fan, 2024].
@@
In recent years, satellite remote sensing data have emerged as promising new data sources for city-wide transportation models, overcoming the weaknesses of localized and mobile ground-based sensors [Sheehan et al., 2023, Gagliardi et al., 2023]. They have demonstrated effectiveness in monitoring both transportation infrastructure conditions and network-wide traffic dynamics. These data often have larger geographical coverage, and provide detailed, consistent information of all individual vehicles detected and road features in city-wide networks. Advances in computer vision have significantly enhanced the information extraction process from these satellite images, including vehicle detection [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016] and road segmentation [Dai et al., 2020, Mei et al., 2021, Kawamura et al., 2025]. Fusing this emerging data source, which is unconstrained by spatial locations, into DODE has great potential to make it more scalable, accurate, and economic.
@@
This study investigates the potential of incorporating high-resolution imagery data collected by flexible sky-view sensing techniques such as satellites into a computation graph-based DODE framework to enhance model scalability and performance in replicating real-world network-wide traffic conditions. The main contributions are summarized as follows:
@@
    We introduce high-resolution satellite imagery as a new observation source for the DODE problem. These images provide class-specific network-wide traffic information, which is not restricted to detector locations, therefore complementing the sparse coverage of localized sensors and providing a scalable way to observe traffic states, especially on links without detectors. They also serve as an additional regularization for DODE to address underdetermined issues when historical OD matrices are not available.
@@
    We integrate a CV pipeline into our DODE framework for density observation extraction. This CV module is compatible with the computational graph-based DODE framework and can be further enhanced to improve detection accuracy. We explicitly model parking and moving vehicles in an agent-based mesoscopic dynamic network loading (DNL) model and use separate dynamic assignment ratio (DAR) matrices for inflow and outflow of curbside parking and moving vehicles to derive gradients for densities.
@@
    We test our model on toy and large-scale networks using both synthetic and real world data. Our experiments reveal the benefit of incorporating satellite-derived density data into DODE and evaluate the sensitivity to sensing errors and sampling frequency. The results show the potential of using this emerging remote sensing data in large-scale applications.
The organization of the remaining paper is as follows: Section 2 summarizes related studies in DODE and the usage of satellite imagery in transportation research. Section 3 introduces the proposed DODE framework integrating multi-source data including satellite imagery information and the general DNL model for density estimation. Section 4 conducts numerical experiments using both real-world multi-source data and synthetic data. Sensitivity analysis is also performed to test the model’s robustness to the data quality of density information extracted from satellite imagery. Section 5 summarizes the findings and provides future research directions.
@@ 2.1 Dynamic OD Demand Estimation
Traditional dynamic origin-destination demand estimation (DODE) has been formulated as an inverse traffic assignment problem, where time-varying OD demand is inferred as inputs to an equilibrium or simulation model to reproduce observed traffic states. We categorize existing DODE studies as model-based, learning-based, and hybrid approaches in this section. Model-based formulations embed a dynamic traffic assignment (DTA) model within a bi-level optimization problem, where the upper level minimizes the discrepancy between observed and simulated traffic states while the lower level propagates traffic flows through a physics-based network flow model subject to behavioral constraints. Learning-based approaches do not rely on explicit traffic assignment models and infer OD flows from observations using data-driven architectures. Hybrid approaches combine both perspectives by embedding learnable components (e.g. neural networks) within DODE pipelines while retaining behavioral and physical consistency.
Recent studies have proposed computational graph-based formulations, enabling end-to-end differentiation and heterogeneous data fusion within DODE pipelines. In this line of work, network loading operators, travel behavior models, and other learnable modules can be encoded as a layered computational graph, enabling efficient gradient-based calibration, which combines model-based and learning-based models within a unified framework. Wu et al. [2018] proposed a multi-layered computational graph representation to structurally model travel demand variables and individual behavior parameters. Ma et al. [2020] and Liu and Qian [2024] formulated DODE problems on a computational graph with a mesoscopic simulation utilizing multi-source system-level data. Lu et al. [2023] proposed a physics-informed framework for joint traffic state and queue profile estimation on layered computational graph with heterogeneous data, and Kim et al. [2024] further integrated transport demand and supply models in the framework. Guarda et al. [2024], Guarda and Qian [2024] developed a computational graph framework to estimate network flow and travel behavior using day-to-day data and further used this framework to estimate traffic conditions for unobserved network locations. Liu et al. [2023b], Liu and Yin [2025] embedded neural networks in implicit layers of an end-to-end learning framework to directly learn travel choice preferences, supply and demand components, and equilibrium state from multi-day observations. Du et al. [2025] developed an end-to-end simulation-based optimization framework to calibrate metro passenger route choice. Zhou et al. [2025] proposed a unified tensor-based computational graph for advanced traffic assignment across multiple modeling layers. These studies demonstrate the potential of this computational-graph framework to improve demand estimation while maintaining consistency with physical and behavioral models and leveraging diverse data sources.
Regardless of the modeling approaches, DODE problems are highly data-demanding. Conventional traffic data used in DODE is collected using either fixed detectors or mobile detectors. Fixed detectors have fixed locations and provide continuous localized traffic conditions at these locations. Some examples are traffic counts collected by loop detectors, automated vehicle identification (AVI) data collected through communication between vehicle tags and barrier devices, and license plate recognition (LPR) data captured by monitoring cameras. Mobile detectors are portable devices attached to a subset of travelers in the system and collect successive traffic conditions based on location changes of these travelers. Some examples are link speed data collected by probe vehicles, vehicle traces of probe vehicles, and human activity trajectories collected by mobile phone location tracking. Table 1 summarizes some representative studies utilizing various data sources within each modeling category.
@@
Dixon and Rilett [2002] automated vehicle identification &\& vehicle traces model-free
@@
Ma et al. [2020] traffic count &\& speed model-based
@@
Cao et al. [2021] automated vehicle identification &\& vehicle traces model-free
@@
One widely acknowledged challenge of DODE is that it is intrinsically an underdetermined problem. Even if links are fully observed, multiple OD patterns may generate the same network conditions, and the relative error of an estimated OD matrix can be unbounded [Yang et al., 1991]. To tackle this challenge, existing literature follows two complementary lines: (1) imposing structural assumptions on the OD matrix and parametric models and (2) leveraging richer and more diverse data sources to add independent constraints.
One representative approach in the first line is low rank based dimension reduction, where temporal or spatial OD patterns are represented using a small number of latent factors. Principal Component Analysis (PCA) [Djukic et al., 2012b, a, Prakash et al., 2017, 2018, Krishnakumari et al., 2020, Qurashi et al., 2022] and compressed sensing techniques [Sanandaji and Varaiya, 2016, Wang et al., 2022] project high-dimensional OD matrices onto a low-dimensional basis, which can be estimated even when historical OD matrices are sparse or partially unavailable. Therefore, low rank-based models can alleviate underdetermination and scalability issues without requiring new data types, by exploiting low rank structure in historical OD patterns. Another category of quasi-dynamic models specifies parametric evolution over time, reduce the number of independent decision variables, and improve scalability on large-scale networks by adding the structural constraints in parametric models [Cascetta et al., 2013]. In addition, utility-based OD estimation and choice-model-driven formulations impose behavioral structure on OD demand [Cantelmo et al., 2018]. By linking OD flows to behavioral models and other demand-side parameters, they effectively shrink the feasible OD space through assumptions on preferences and responses to various attributes.
The second line of work retains a high-dimensional OD representation and addresses underdetermined nature by incorporating rich, independent observations as regularization. Prior studies have incorporated traffic counts and travel times [Ma et al., 2020], probe trajectories [Dixon and Rilett, 2002, Cao et al., 2021, Carrese et al., 2017], and other mobility datasets [Liu and Qian, 2024, Wang et al., 2013, Lu et al., 2024] to better constrain optimization problem and estimate network flow patterns. These heterogeneous data sources are complementary but each has distinct characteristics and limitations. Fixed detectors can be expensive and spatially sparse, and sensitive to sensor placement [Hu and Fan, 2024]. Additionally, maintaining installed sensors and planning new locations can be costly, especially in large-scale networks, limiting scalability and sustainability. Mobile detectors are more flexible but rely on penetration rates and may suffer from biased sampling, incomplete coverage, and privacy issues [He et al., 2024]. Moreover, both sensing categories can exhibit inconsistent data quality across time and space, potentially leading to contradictory measurements for the same link. To overcome these limitations, researchers have explored alternative observations that capture traffic states more directly. For example, Frederix et al. [2010] combined additional link-level density information into DODE, outperforming traditional estimation approaches using only traffic flow because density observations are effective in distinguishing congested and uncongested traffic states. In this context, emerging remote sensing (e.g., satellite imagery) provides network-wide snapshots of traffic states (density) with broad spatial coverage but limited temporal frequency, making it a promising complementary data source for DODE. This study follows this line of literature and explores the potential of incorporating satellite image-derived densities into DODE as an additional data source.
@@ 2.2 Satellite imagery in transportation research
Advances in computer vision and sensing have unlocked the potential of remote sensing data for enhancing city-wide infrastructure monitoring [Gagliardi et al., 2023] and traffic condition monitoring [Sheehan et al., 2023]. High-resolution imagery data captures consistent, city-wide traffic conditions across all links at the same time, overcoming the limitations of ground-based sensors. Some pilot studies have demonstrated the effectiveness of traffic monitoring using satellite imagery data [McCord et al., 2003, Larsen et al., 2009, 2013, Kaack et al., 2019], and advanced computer vision algorithms support the extraction of fine-grained information from satellite images [Kawamura et al., 2025]. Additionally, Unmanned Aerial Vehicles (UAV) can be another similar but more flexible approach to collect precise imagery or video-based traffic data in small sub-regions of a city, due to their rapid speed, cost-effectiveness, and extensive field-of-view [Balamuralidhar et al., 2021, Butilă and Boboc, 2022, Bai and Feng, 2025, Fonod et al., 2025, Espadaler-Clapés et al., 2025, Xiong et al., 2025].
@@
Some pilot studies have explored the utilization of these imagery data for transportation research, including road network extraction [Dai et al., 2020, Mei et al., 2021], vehicle detection and classification [Leitloff et al., 2010, Chen et al., 2014, Cao et al., 2016]. Some studies have focused on traffic density estimation based on imagery or video collected from satellites [Kopsiaftis and Karantzalos, 2015, Sakai et al., 2019]. Reksten and Salberg [2021] proposed a deep learning-based framework to estimate annual average daily traffic (AADT) using very-high-resolution optical remote-sensing imagery with spatially limited ground-based measurements. Ganji et al. [2022] used Google aerial images together with sparse data from traffic counts to predict traffic volumes. Chen et al. [2021] used traffic density detection data to examine changes in spatio-temporal traffic patterns during the COVID-19 pandemic. He et al. [2021] integrated satellite imagery and GPS trajectories to generate high-resolution accident risk maps. Yu et al. [2024] integrated high-resolution satellite imagery with geo-spatial tabular and crash data to analyze factors contributing to freight truck-related crashes. Horsler et al. [2023] incorporated information from satellite imagery to predict regional road transport emissions, highlighting the potential in environmental impact assessments. Our initial work demonstrates the potential of incorporating satellite imagery into mesoscopic network model calibration but the information extraction from imagery data is not vehicle class specific and the parking information is not considered, leading to limited accuracy in large-scale network applications [Liu et al., 2024].
@@
To the best of our knowledge, none of the existing studies have incorporated detailed traffic information of both moving and on-street parking vehicles by class extracted from satellite images directly into calibrating dynamic network models and performing DODE. Incorporating satellite imagery has the potential to fill gaps in areas where ground-based sensors are not available, providing a more complete picture of consistent network conditions to enhance DODE performance.
Figure 1: Model Overview